OpenStudy (anonymous):

simplify sqrt(32)-sqrt(8). Please help.

OpenStudy (anonymous):

We can write sqrt32=sqrt(2*2*2*2*2)=sqrt(2^2*2^2*2)=(sqrt(2^2)*sqr(2^2)*sqrt2)=2*2*sqrt2=4sqrt2 and sqrt8=sqrt(2*2*2)=(sqrt(2^2)*sqrt2)=2sqrt2 so...we can rewrite our equation as..4sqrt2-2sqrt2=2*2sqrt2-2sqrt2.. taking 2sqrt2 as common out..we will get 2sqrt2(2-1)=2sqrt2 HOPE THIS WILL HELP YOU

OpenStudy (anonymous):

you can also try it like...sqrt(8*4)-sqrt8=sqrt8*sqrt4-sqrt8=sqrt8(sqrt4-1)=2sqrt2(2-1)=2sqrt2